# v-Hub: A Visual-Centric Humor Understanding Benchmark for Video LLMs

### v1

@@ 6 Conclusion
We have introduced v-Hub, a visual-centric humor understanding benchmark. v-Hub is designed to assess and diagnose the capability of MLLMs for video humor understanding. It contains a collection of funny videos collected from two complementary domains. Each clip is annotated with captions, descriptions, explanations, etc., supporting evaluation tasks such as caption matching and humor explanation. To broaden the applicability of v-Hub, we further construct an open-ended task, contributing to a comprehensive evaluation of MMLMs for video understanding. We evaluated a diverse range of MMLMs, spanning open-sourced and proprietary domains and covering specialized video LLMs and versatile OmniLLMs. Our findings reveal that current MLLMs heavily rely on linguistic cues for humor understanding, but are weak in deriving nuanced visual cues for understanding sophisticated video humor. Moreover, we empirically find that including audio is helpful for humor understanding, highlighting the informativeness of sound and the promise of incorporating rich modalities for complex video reasoning tasks.
Ethics statement
Our work follows widely recognized ethical principles in computing research, including the ACM Code of Ethics. In developing our benchmark, we considered the following aspects:
  • Contribute to Society and to Human Well-being: Our dataset is intended to advance multimodal AI research on humor understanding, a capability that has broad applications in safe content moderation, assistive technologies and cross-cultural communication. We employed three Human Intelligence Tasks (HITs) to gather data, and we fully considered cultural and linguistic diversity in humor to minimize potential negative impacts as much as possible, such as reinforcing harmful stereotypes, exposing sensitive content, or infringing on personal safety and privacy. To ensure accessibility and inclusivity, the dataset will be made broadly available for non-commercial research purposes.
  • Uphold High Standards of Scientific Excellence: We consistently adhere to principles of transparency and rigor throughout dataset curation and analysis. We meticulously document all preprocessing, annotation and evaluation procedures, and we will publicly release the relevant code to ensure independent verification and reproducibility. Furthermore, we provided fair compensation to all annotation personnel, ensuring their hourly wages exceeded the local minimum wage. Finally, we did not fabricate, falsify, or misrepresent data. Beyond the data annotators, no other human subjects were directly involved, and no personally identifiable information was used. Therefore, no additional ethics approval is required.
  • Avoid Harm: We carefully considered potential risks that could arise from constructing and releasing a humor video benchmark. To mitigate negative consequences, we implemented a multi-stage screening mechanism to exclude humorous content featuring violence, discrimination, or relying on stereotypes targeting vulnerable groups. Simultaneously, we adopted a three-person collaborative annotation scheme, ensuring that each data entry underwent three rounds of annotation. We explicitly document the limitations of the dataset to minimize unintended harm arising from cross-cultural or linguistic misunderstandings. Furthermore, we analyzed potential downstream risks, such as misuse for generating harmful or offensive humor, and explicitly warn against such applications in the dataset license.
  • Be Honest, Trustworthy and Transparent: We transparently disclose the characteristics, strengths, and limitations of the dataset. While this dataset thoroughly considers cultural and linguistic diversity in humor and encompasses diverse humor scenarios, it cannot cover all dimensions of global humor cultures. Furthermore, Charlie Chaplin’s Silent Films constitute a significant portion of the dataset, inevitably introducing potential cultural bias. We confirm that there are no conflicts of interest that may compromise the independence of our research, and all funding sources are clearly acknowledged. At the same time, we guarantee that we do not misrepresent related work, nor do we claim capabilities beyond what our benchmark enables.
  • Be Fair and Take Action not to Discriminate: We strived for fairness in both dataset curation and evaluation. And we made efforts to avoid humor that demeans particular groups. We also emphasize that the dataset should not be used to develop systems that discriminate, disenfranchise, or oppress individuals.
  • Respect the Work Required to Produce New Ideas and Artefacts: We credit the creators of ideas, inventions, work, and artefacts, and respect copyrights and property. All videos are sourced from publicly available materials with appropriate licenses. Where possible, we provide attribution to content creators and respect cultural heritage by excluding sensitive or protected media. Where possible, we will provide attribution to content creators and respect their work by excluding sensitive or protected media content.
  • Respect Privacy: Our work did not use private or personally identifiable data. All videos have been anonymized or utilize publicly available Charlie Chaplin silent films, carefully mitigating the risk of re-identification. Furthermore, the dataset is restricted to legitimate academic research under the dataset license.
  • Honour Confidentiality: Our work did not involve any confidential or proprietary information. Reviewers and collaborators were only provided with materials approved for release. We commit to maintaining confidentiality in peer review and in handling sensitive communications.
Appendix A Additional related work
A.1 From LLMs to Video LLMs
Large language models have demonstrated outstanding capabilities in many domains, including natural language processing, coding, math, and reasoning, ushering in new breakthroughs for video understanding technology. Video LLMs integrate visual encoders with LLMs, leading to a unified model to reason across video and language in the same language space (Wang et al., 2024a; Liu et al., 2024; Lin et al., 2023). Early video LLMs employ pre-trained image encoder and video encoder to encode only video frames  (Zhang et al., 2023; Maaz et al., 2023; Li et al., 2023; Lin et al., 2023). Recent works augment video LLMs with an audio encoder to align visual, auditory, and textual modalities in the same language space. Moreover, the audio encoder is supposed to capture diverse environmental sound apart from human speech since sound has been shown to contain amounts of commonsense knowledge  (Cheng et al., 2024; Xu et al., 2025).
Appendix B Crowdworking Details
B.1 Processing and filtering
@@ B.2 Annotation
Qualification.The recruited crowdworkers were mainly from China, all possessing at least an undergraduate education and with English background. Before formal annotation began, we conducted training sessions and a qualification review. During the qualification stage, crowdworkers were required to annotate 15 video samples. We manually reviewed their results and assigned scores based on the annotation guidelines. Ultimately, we selected eight qualified annotators.
@@
Appendix C Data Statistics
(a) Distribution of video lengths
(b) Distribution of humor types
Figure 7: Data statistics: video length and humor type distributions.
Duration.
Diversity.
To show our dataset contains a variety of humor, we follow Buijzen & Valkenburg (2004) to categorize humor into five categories (see Figure 7(b)): Slapstick humor, Clownish humor, Surprise, Irony, Misunderstanding and Others (e.g. Parody, Miscellaneous, Satire).
Visual-centric.
Our dataset is predominantly visual centric, with 99% of videos relying primarily on visual cues. Specifically, per the four groups defined in Section 2.4: Visual, Visual+Text, Visual+Audio, and Visual+Autio+Text, 600 videos (63%) fall into the category Visual, meaning humor is entirely derived from facial expressions, object interactions, or visual effects without reliance on text or sound effects. 214 videos (22%) integrate audio, indicating that while humor remains visually driven, auditory elements such as background music or sound effects enhance the comedic impact. 92 videos (9%) are classified as Visual+Text indicating that video humor is extended through additional visual text. 46 videos (5%) combines the three modalities to deliver humor. The remaining 1% includes videos, in which speech plays a minor role but does not dominate the expression of humor.
Figure 8: Video distribution over four groups.
Appendix D Additional Experimental Notes
D.1 DETAILS OF GENERATE OPEN-ENDED QA PAIRS
@@
D.2 Details of evaluation methods
BERTScore.
It evaluates the semantic similarity between generated answers and ground truth (GT) references by leveraging contextual embeddings from pre-trained language models (Zhang* et al., 2020). Instead of relying on exact lexical overlap, BERTScore computes similarity at the token level in embedding space and aggregates these scores into three metrics:
  • Recall measures how well the GT tokens are semantically captured by the generated answer.
  • Precision measures how relevant the generated tokens are with respect to the GT reference.
  • F1 Score is the harmonic mean of precision and recall, providing a balanced assessment of semantic alignment.
The inclusion of BERTScore allows us to evaluate whether model-generated answers semantically align with human references, thereby assessing the adequacy and quality of humor comprehension.
@@ D.3 Baseline Models
Qwen-2.5-Omni 7B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
To evaluate multimodal large language models’ ability to understand video humor, we selected state-of-the-art models representing three distinct input modalities, as summarized in Table 6. Specifically, we include multimodal LLMs that process raw visual frames and text, and omni LLMs that integrate both text, video and audio signals. This set covers both public models (e.g., Qwen2.5-VL-72B, Intern3.5-VL) and proprietary models (e.g., Gemini-2.5-flash, GPT-4o), offering a broad perspective on current approaches. Each model is evaluated under all input conditions it can handle (see Section 4.1): for instance, omni-modal models can participate in the Text-Only, Video-Only, and Video+Audio groups, whereas multimodal models are tested exclusively with textual input and raw visual frames. This setup allows us to isolate how each model category—multimodal and omni-modal—contributes to humor understanding across diverse input modalities.
Appendix E Additional Experimental Results
Models Open-ended QA
SentBERT METEOR Precision Recall F1 Score
Qwen2.5-VL-72B 0.792 0.622 0.744 0.740 0.738
Qwen2.5-72B 0.710 0.636 0.680 0.723 0.700
Qwen2.5-Omni-7B 0.740 0.555 0.703 0.703 0.698
Qwen2.5-7B 0.692 0.539 0.657 0.688 0.667
MLLMs vs. their base LLMs.
MLLMs are usually derived from a pre-trained base LLM by adding a visual encoder or multimodal modules. For instance, Qwen2.5-VL-72B extends Qwen2.5-72B, and Qwen2.5-Omni extends Qwen2.5-7B (see Table 7). In the Text-Only setup, Qwen2.5-Omni surpasses Qwen2.5-7B with a SentBERT score of 0.740 (vs. 0.692) and an F1 score of 0.698 (vs. 0.667) on Open-ended QA task, suggesting that multimodal training can confer advantages even when only textual descriptions are available—possibly because the model has learned richer contextual associations during training. Please refer to Table 11 for more details on humor explanation and caption matching tasks.
Table 8: The impact of requiring background knowledge support on video humor understanding.
@@ Appendix E Additional Experimental Results
SentBERT METEOR F1 Score Accuracy SentBERT METEOR F1 Score
Gemini-2.5-flash 0.503 0.211 0.568 0.633 0.437 0.268 0.550
video-SALMONN-2 0.273 0.153 0.513 0.266 0.306 0.159 0.520
MiniCPM2.6-o 0.404 0.166 0.532 0.378 0.318 0.103 0.442
Qwen-2.5-Omni 0.404 0.157 0.531 0.571 0.384 0.105 0.486
Background knowledge does not necessarily improve video humor understanding.
The results in Table 2 and Table 8 show that there is no significant difference between the mean scores and the scores for videos that need background knowledge to perceive humor under Video+Audio setting. For example, under the Video+Audio setting, the Gemini-2.5-flash attains an average F1 score of on the Explanation task for Background-Dependent videos 0.568, which is statistically similar to its F1 score of 0.550 on the full dataset. This suggests that the language-model component of MLLMs already encodes most of the cultural background knowledge necessary for humor comprehension, meaning that the absence of explicit background knowledge in the input does not significantly degrade their performance. MLLMs do not show a significant disadvantage in understanding videos that require background knowledge compared to those that do not, potentially because video humor rarely relies on specific background knowledge, making it universally understandable.
@@
The results in Table 9 indicate that current MLLMs rely heavily on linguistic cues to generate reasonable explanations, and struggle to effectively extract semantic information from raw visual or auditory signals. For example, Qwen2.5-VL-72B attains a preference score of 0.687 under Text-Only, significantly outperforming its Video-Only score of 0.423. Furthermore, although closed-source models demonstrate greater robustness under multimodal inputs, they still struggle to align visual and audio cues to enhance humor comprehension. For instance, Gemini-2.5-flash achieves 0.546 (Video-Only) and 0.566 (Video+Sound).
Appendix F The Use of Large Language Models
In this work, we used LLMs as assistive tools in several stages of the research:
  • Dataset construction. We initially employed GPT-4o to assist in generating candidate QA pairs and humor categories from video content. All outputs were subsequently reviewed and revised by human annotators to ensure correctness and quality.
  • Writing support. LLMs were used to write and polish some sentences in the paper for readability.
@@ Appendix F The Use of Large Language Models
Qwen2.5-VL-72B [Uncaptioned image] MLLM 0.687 0.423
@@
Table 10: Model performance on Humor Explanation, Caption Matching, and Open-ended QA.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
Text-Only
Gemini-2.5-flash 0.550 0.616 0.580 0.611 0.710 0.744 0.723
video-SALMONN-2 0.570 0.612 0.589 0.367 0.650 0.636 0.639
MiniCPM2.6-o 0.519 0.614 0.562 0.518 0.460 0.666 0.543
Qwen-2.5-Omini 0.540 0.605 0.570 0.644 0.703 0.703 0.698
Qwen-2.5-VL-72B 0.548 0.615 0.578 0.719 0.744 0.740 0.738
Intern3.5-VL 0.548 0.618 0.580 0.632 0.688 0.701 0.689
GPT-4o 0.546 0.626 0.581 0.767 0.698 0.746 0.718
Video-Only
Gemini-2.5-flash 0.522 0.583 0.550 0.583 0.547 0.570 0.556
video-SALMONN-2 0.487 0.525 0.504 0.259 0.549 0.509 0.525
MiniCPM2.6-o 0.500 0.543 0.520 0.364 0.397 0.522 0.452
Qwen-2.5-Omini 0.503 0.543 0.521 0.55 0.538 0.457 0.488
Qwen-2.5-VL-72B 0.526 0.571 0.547 0.673 0.567 0.541 0.550
Intern3.5-VL 0.522 0.568 0.543 0.640 0.544 0.545 0.542
GPT-4o 0.515 0.585 0.547 0.667 0.538 0.579 0.556
Video+Sound
Gemini-2.5-flash 0.521 0.583 0.550 0.588 0.543 0.570 0.554
video-SALMONN-2 0.487 0.529 0.506 0.255 0.556 0.526 0.538
MiniCPM2.6-o 0.501 0.553 0.523 0.404 0.494 0.541 0.513
Qwen-2.5-Omini 0.509 0.558 0.531 0.623 0.565 0.507 0.529
@@
SentBERT METEOR Precision Recall F1 Score Accuracy
Qwen2.5-VL-72B 0.553 0.250 0.548 0.615 0.553 0.719
Qwen2.5-72B 0.546 0.245 0.552 0.615 0.581 0.646
Qwen2.5-Omni-7B 0.547 0.232 0.605 0.570 0.547 0.644
Qwen2.5-7B 0.568 0.241 0.540 0.613 0.573 0.522
Table 12: The impact of requiring background knowledge support on video humor understanding.
Models Explanation Matching Open-ended QA
Precision Recall F1 Score Accuracy Precision Recall F1 Score
Gemini-2.5-flash 0.545 0.595 0.568 0.633 0.540 0.566 0.550
Qwen-2.5-Omni 0.517 0.547 0.531 0.571 0.533 0.460 0.486
MiniCPM2.6-o 0.516 0.550 0.532 0.378 0.396 0.521 0.442
video-SALMONN 2 0.501 0.527 0.513 0.266 0.542 0.507 0.520
Table 13: The impact of background knowledge on video humor understanding.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
@@
video-SALMONN-2 0.561 0.568 0.563 0.331 0.554 0.525 0.535
MiniCPM-2.6-o 0.533 0.585 0.557 0.445 0.508 0.553 0.520
Qwen-2.5-Omni 0.536 0.582 0.557 0.667 0.583 0.540 0.556
@@
video-SALMONN-2 0.499 0.533 0.515 0.261 0.545 0.517 0.528
MiniCPM-2.6-o 0.521 0.563 0.540 0.412 0.490 0.537 0.509
Qwen-2.5-Omni 0.462 0.181 0.545 0.630 0.445 0.158 0.526
Table 14: The impact of video era on video humor understanding.
Explanation Matching Open-ended QA
Models Precision Recall F1 Score Accuracy Precision Recall F1 Score
Charlie Chaplin’s Silent Films
Gemini-2.5-flash 0.517 0.570 0.541 0.562 0.529 0.567 0.545
video-SALMONN-2 0.492 0.528 0.509 0.165 0.521 0.511 0.513
MiniCPM2.6-o 0.492 0.527 0.508 0.307 0.431 0.535 0.470
Qwen-2.5-Omni 0.495 0.528 0.510 0.494 0.515 0.482 0.493
User-Generated Funny Video
Gemini-2.5-flash 0.525 0.588 0.553 0.592 0.553 0.572 0.560
video-SALMONN-2 0.485 0.523 0.503 0.296 0.559 0.509 0.529
MiniCPM2.6-o 0.503 0.549 0.524 0.385 0.405 0.518 0.446
Qwen-2.5-Omni 0.506 0.549 0.525 0.571 0.547 0.448 0.487
Appendix G Prompts
We list our prompt in Tables˜15, 16, 19, 20, 22, 21, 18, 17, 25, 24 and 23
Table 15: Prompt for writing captions of videos.
Requirements:
Please ensure it is related to the video content.
Output format:
Caption: <caption>
Video description: {video_description}
Descriptive captions: {descriptive_captions}

### v2

@@ 6 Conclusion
We have introduced v-Hub, a video humor understanding benchmark from vision and sound. v-Hub is designed to assess and diagnose the capability of MLLMs for video humor understanding. Each video is annotated with captions, descriptions, explanations, etc., supporting evaluation tasks such as caption matching, humor explanation, and open-ended QA. We evaluated a diverse range of MLLMs, spanning open-sourced and proprietary domains and covering specialized video LLMs and versatile OmniLLMs. Our findings reveal that current MLLMs heavily rely on linguistic cues for humor understanding, but are weak in deriving nuanced visual cues for understanding sophisticated video humor. Moreover, we empirically find that including environmental sound helps with humor understanding, highlighting the informativeness of sound and the promise of incorporating rich modalities for complex video reasoning tasks.
Limitations
We acknowledge two limitations of this work: (1) v-Hub contains humorous videos from diverse cultural backgrounds. Although culture-level categorization would be more fine-grained, assigning a single culture label to visual-centric humorous videos is often ambiguous and subjective, due to the prevalence of hybrid and globally shared visual humor. We therefore adopt region-level categorization, which is proposed by House et al. (2004) and Ronen and Shenkar (2013), as a more stable, weakly supervised proxy for cultural background, while acknowledging that a geographical region may encompass multiple cultures. (2) Second, all data annotations were created manually in English. However, the ability of LLMs to understand humor may vary across languages, as different languages interpret and express the same visual content in distinct ways. This points to a promising direction for future research.
Ethical Considerations
We firmly adhere to the ACL Code of Ethics in the performance of this work and the methods involved. We respect intellectual property and privacy rights and restrict the dataset to non-commercial academic use (see more details in Appendix A.3). Additionally, we carefully screened user-generated content to ensure the benchmark is safe for the research community and minimizes privacy risks for private individuals. Moreover, LLMs may produce offensive and incorrect statements, explicitly warn against the misuse of this dataset for generating malicious or mocking content, advocating instead for the development of safe and empathetic AI systems. This work is released with the intent for research purposes only.
Acknowledgements
Yanpeng Zhao acknowledges the support of the National Natural Science Foundation of China (12574467). We would like to thank Hengli Li for their helpful suggestions. We are also grateful to all annotators who contributed to the construction and verification of v-Hub. Their help and efforts were essential to ensuring the quality and reliability of the benchmark.
Use of LLMs.
In this work, we used LLMs as assistive tools in the following stages:
  • Dataset Construction. We initially employed GPT-4o (Hurst et al., 2024) to assist in generating candidate QA pairs and humor categories from video content. All outputs were subsequently reviewed and revised by human annotators to ensure correctness and quality.
  • Writing Support. ChatGPT was used to write and polish some sentences in Section 4 for readability.
Appendix A Crowdworking Details
A.1 Processing and Filtering
@@ A.2 Annotation
Qualification. The recruited crowdworkers were mainly from China, all possessing at least an undergraduate education and with English background. Before formal annotation began, we conducted training sessions and a qualification review. During the qualification stage, crowdworkers were required to annotate 15 video samples. We manually reviewed their results and assigned scores based on the annotation guidelines. Ultimately, we selected eight qualified annotators. And we provided fair compensation to all crowdworkers, ensuring their hourly wages exceeded the local minimum wage.
Figure 5: Examples of visual-centric humor understanding, where ‘text’ refers to visual text.
@@
A.3 Copyright & License
We respect the copyright of each video. We have emailed Charlie Chaplin’s copyright holders regarding copyright issues related to Chaplin clips, and v-HUB is only used for academic research. Commercial use in any form is prohibited. The copyright of all videos belongs to the video owners, and we will remove the videos upon their request. Without prior approval, you cannot distribute, publish, copy, disseminate, or modify v-HUB in whole or in part. You must strictly comply with the above restrictions.
@@ A.3 Copyright & License
Appendix B Additional Experimental Notes
B.1 Details of Generate Open-ended QA Pairs
@@
B.2 Details of Evaluation Methods
We used the 3.9.1 NLTK, 0.3.13 bert-score, and 5.1.1 sentence-transformers package to calculate the metrics.
AutoDQ.
It evaluates the presence of humor-related events in the generated explanations (Wang et al., 2024a). It extracts key events from the model’s output and compares them to ground truth (GT) annotations using entailment analysis. It provides three metrics: recall, precision, and F1 score, defined as:
  • Recall measures the percentage of GT events entailed by the model-generated events.
  • Precision measures the percentage of model-generated events that are entailed by GT events.
  • F1 Score is the harmonic mean of precision and recall, serving as a balanced indicator that jointly reflects coverage and correctness.
The inclusion of AutoDQ allows us to evaluate factual correctness and event coverage, checking whether the explanations cover all the humorous points in the video content.
@@ B.3 Baseline Models
Qwen2.5-Omni 7B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
To evaluate multimodal large language models’ ability to understand video humor, we selected state-of-the-art models representing three distinct input modalities, as summarized in Table 7. Specifically, we include multimodal LLMs that process raw visual frames and text, and omni LLMs that integrate both text, video and audio signals. This set covers both public models (e.g., Qwen2.5-VL, Intern3.5-VL) and proprietary models (e.g., Gemini-2.5-Flash, GPT-4o), offering a broad perspective on current approaches. Each model is evaluated under all input conditions it can handle (see Section 4.1): for instance, omni-modal models can participate in the Text-Only, Video-Only, and Video+Audio groups, whereas multimodal models are tested exclusively with textual input and raw visual frames. This setup allows us to isolate how each model category—multimodal and omni-modal—contributes to humor understanding across diverse input modalities.
Appendix C Additional Experimental Results
Table 8: The impact of requiring background knowledge support on video humor understanding in the Video-Only setting.
@@ Appendix C Additional Experimental Results
SentBERT METEOR AutoDQ Accuracy SentBERT METEOR
Background-Dependent videos
Gemini-2.5-Flash 0.500 0.211 0.198 0.628 0.433 0.266
video-SALMONN-2 0.271 0.153 0.028 0.257 0.307 0.160
MiniCPM2.6-o 0.402 0.166 0.093 0.374 0.328 0.120
Qwen2.5-Omni 0.401 0.158 0.137 0.559 0.382 0.105
Full dataset
Gemini-2.5-Flash 0.459 0.199 0.175 0.580 0.424 0.270
video-SALMONN-2 0.269 0.150 0.052 0.243 0.317 0.169
MiniCPM2.6-o 0.381 0.165 0.112 0.362 0.369 0.186
Qwen2.5-Omni 0.384 0.159 0.144 0.553 0.382 0.121
Comparable performance on videos with and without background knowledge requirements.
The results in Table 8 show that model performance on Background-Dependent videos is largely comparable to that on the full dataset in the Video-Only setting. For example, video-SALMONN-2 attains an average SentBERT of 0.271 on the humor explanation task for background-dependent videos, which is statistically similar to its SentBERT of 0.269 on the full dataset. This suggests that the language-model component of MLLMs already encodes most of the cultural background knowledge necessary for humor comprehension, meaning that the absence of explicit background knowledge in the input does not significantly degrade their performance. MLLMs show a comparable performance in understanding videos that require background knowledge compared to those that do not, potentially because video humor rarely relies on specific background knowledge, making it universally understandable.
@@
The results in Table 9 indicate that current MLLMs rely heavily on linguistic cues to generate reasonable explanations, and struggle to effectively extract semantic information from raw visual or auditory signals. For example, Qwen2.5-VL attains a preference score of 0.687 under Text-Only, significantly outperforming its Video-Only score of 0.423. Furthermore, although closed-source models demonstrate greater robustness under multimodal inputs, they still struggle to align visual and audio cues to enhance humor comprehension. For instance, Gemini-2.5-Flash achieves 0.546 (Video-Only) and 0.566 (Video+Audio).
Appendix D Case Study
We present our case studies in Figure 9.
@@ Appendix D Case Study
Qwen2.5-VL [Uncaptioned image] MLLM 0.687 0.423
@@
(a) case study 1
(b) case study 2
(c) case study 3
Figure 9: Three case studies illustrating our findings.
Table 10: The impact of visual text on video humor understanding in the Video+Audio setting.
Sound contributing to humor Sound not contributing to humor
Explanation Open-ended QA Explanation Open-ended QA
Models SentBERT METEOR BERTScore Precision Recall F1 Score BERTScore SentBERT METEOR BERTScore Precision Recall F1 Score BERTScore
w/ visual text
Gemini-2.5-Flash 0.458 0.199 0.541 0.157 0.210 0.174 0.552 0.472 0.206 0.547 0.185 0.174 0.177 0.550
video-SALMONN-2 0.258 0.164 0.481 0.169 0.042 0.067 0.532 0.261 0.160 0.509 0.077 0.045 0.057 0.540
MiniCPM2.6-o 0.416 0.170 0.516 0.105 0.111 0.108 0.527 0.434 0.182 0.522 0.143 0.137 0.151 0.536
Qwen2.5-Omni 0.441 0.170 0.522 0.148 0.133 0.140 0.521 0.430 0.176 0.522 0.159 0.133 0.145 0.538
w/o visual text
Gemini-2.5-Flash 0.482 0.202 0.550 0.165 0.195 0.179 0.539 0.448 0.195 0.543 0.034 0.051 0.041 0.552
video-SALMONN-2 0.276 0.185 0.496 0.124 0.045 0.064 0.529 0.285 0.170 0.504 0.098 0.034 0.050 0.541
MiniCPM2.6-o 0.407 0.182 0.525 0.130 0.097 0.103 0.515 0.364 0.167 0.510 0.060 0.048 0.046 0.514
Qwen2.5-Omni 0.425 0.181 0.538 0.116 0.128 0.122 0.509 0.413 0.173 0.522 0.062 0.045 0.050 0.529
@@
SentBERT METEOR BERTScore Precision Recall F1 Score Accuracy
Qwen2.5-VL 0.543 0.250 0.573 0.323 0.364 0.342 0.719
Qwen2.5-72B 0.536 0.245 0.576 0.318 0.374 0.344 0.661
Qwen2.5-Omni 0.536 0.233 0.565 0.303 0.331 0.316 0.644
Qwen2.5-7B 0.560 0.240 0.567 0.293 0.369 0.324 0.542
Table 12: Model performance on Humor Explanation.
Explanation Open-ended QA
Models BERTScore Precision Recall F1 Score BERTScore
Text-Only
Gemini-2.5-Flash 0.575 0.307 0.385 0.342 0.712
video-SALMONN-2 0.586 0.290 0.350 0.317 0.639
MiniCPM2.6-o 0.558 0.307 0.345 0.325 0.536
Qwen2.5-Omni 0.565 0.303 0.331 0.316 0.687
Qwen2.5-VL 0.573 0.323 0.364 0.342 0.730
Intern3.5-VL 0.576 0.320 0.382 0.348 0.685
GPT-4o 0.574 0.339 0.417 0.374 0.699
Video-Only
Gemini-2.5-Flash 0.546 0.154 0.206 0.176 0.549
video-SALMONN-2 0.497 0.087 0.042 0.052 0.525
MiniCPM2.6-o 0.516 0.116 0.109 0.112 0.469
Qwen2.5-Omni 0.518 0.169 0.126 0.144 0.497
Qwen2.5-VL 0.542 0.150 0.152 0.150 0.546
Intern3.5-VL 0.537 0.126 0.116 0.125 0.540
GPT-4o 0.536 0.214 0.205 0.206 0.544
Video+Sound
Gemini-2.5-Flash 0.546 0.153 0.200 0.173 0.549
video-SALMONN-2 0.499 0.126 0.047 0.066 0.537
MiniCPM2.6-o 0.519 0.122 0.062 0.110 0.514
Qwen2.5-Omni 0.525 0.137 0.116 0.120 0.529
Table 13: The impact of requiring background knowledge support on video humor understanding in the Video-Only setting.
Models Explanation Open-ended QA
BERTScore Precision Recall F1 Score BERTScore
Background-Dependent videos
Gemini-2.5-Flash 0.566 0.193 0.203 0.198 0.549
video-SALMONN-2 0.512 0.051 0.019 0.028 0.488
MiniCPM2.6-o 0.531 0.109 0.079 0.093 0.446
Qwen2.5-Omni 0.530 0.184 0.109 0.137 0.519
Full dataset
Gemini-2.5-Flash 0.546 0.154 0.206 0.176 0.549
video-SALMONN-2 0.497 0.087 0.042 0.052 0.525
MiniCPM2.6-o 0.516 0.116 0.109 0.112 0.469
Qwen2.5-Omni 0.518 0.169 0.126 0.144 0.497
Table 14: The impact of background knowledge on video humor understanding in the Video+Audio setting.
Explanation Open-ended QA
Models BERTScore Precision Recall F1 Score BERTScore
@@
video-SALMONN-2 0.562 0.117 0.107 0.114 0.536
MiniCPM2.6-o 0.555 0.192 0.194 0.193 0.520
Qwen2.5-Omni 0.557 0.195 0.160 0.176 0.555
@@
video-SALMONN-2 0.514 0.084 0.014 0.025 0.528
MiniCPM2.6-o 0.538 0.132 0.103 0.115 0.509
Qwen2.5-Omni 0.544 0.142 0.114 0.127 0.525
Table 15: The impact of video era on video humor understanding in the Video-Only setting.
Models Explanation Open-ended QA
BERTScore Precision Recall F1 Score BERTScore
Last-Century Charlie Chaplin’s Silent Films
Gemini-2.5-Flash 0.541 0.118 0.145 0.130 0.545
video-SALMONN-2 0.509 0.035 0.007 0.012 0.513
MiniCPM2.6-o 0.508 0.116 0.083 0.097 0.470
Qwen2.5-Omni 0.510 0.153 0.070 0.096 0.493
Contemporary User-Generated Funny Video
Gemini-2.5-Flash 0.547 0.173 0.222 0.194 0.550
video-SALMONN-2 0.511 0.103 0.052 0.061 0.530
MiniCPM2.6-o 0.518 0.113 0.124 0.118 0.470
Qwen2.5-Omni 0.520 0.173 0.130 0.166 0.498
Appendix E Prompts
We list our prompt in Tables˜22, 16, 19, 20, 23, 21, 18, 17, 26, 25 and 24